OpenStudy (bass53):

If a ramp to the top of a 30 meter building at a construction site has a mechanical advantage of 4, how long must the ramp be? @Abhisar @ganeshie8

OpenStudy (anonymous):

If we assume that the mechanical advantage is 4 and the 30 meters is the height of the ramp (same as the length of the building), the length can be calculated from the mechanical advantage equation m = l/h , so 4 = l/30, so l = 4 * 30 = 120. The ramp should be 120 meters long.

OpenStudy (abhisar):

You probably didn't noticed but you posted your physics question in the Biology section. Next time check if u r in ryt section ;)

OpenStudy (bass53):

ive already got it thanks tho :)